OpenStudy (anonymous):

Describe the usefulness of the conjugate and its effect on other complex numbers

OpenStudy (anonymous):

Conjugation of a complex number describes an axial symmetry of the complex plane. To conjugate a complex number, reflect its position through the real axis. This geometric significance is used a lot; e.g., http://www.amazon.com/Abels-Theorem-Problems-Solutions-International/dp/1402021860

OpenStudy (anonymous):

@Samie509 did u get it

OpenStudy (anonymous):

Not quite but im getting there # thank you :)

OpenStudy (unklerhaukus):

when a complex number is multiplied by its complex conjugate you get the numbers measure or modulus \[\left|a+ib\right|=\sqrt{(a+ib)\overline{(a+ib)}}=\sqrt{(a+ib)(a-ib)}=\sqrt{a^2+b^2}\]

OpenStudy (anonymous):

Thank yOu wt is its usefullness tho :)